# The Faiss Library

### v2

@@ 6.2 Filtered search
This is sufficient to store enumerated types (e.g. . country codes, music genres, license types, etc.), dates (as days since some origin), version numbers, etc. However, it is insufficient for more complex metadata. In the example use case below, we use the available bits to implement more complex filtering.
@@
In the filtered search track of the BigANN 2023 competition [big,], each query and database vector is associated with a few terms from a fixed vocabulary of size $v$ (for the queries there are only 1 or 2 words). The filtering consists in considering only the database vectors that include all the query terms. . This metadata is given as a sparse matrix $M_{\mathrm{meta}}\in\{0,1\}^{N\times v}$.
@@
$\{w_{1},w_{2}\}\subset W_{i}\Rightarrow\neg s_{i}\wedge s_{\mathrm{q}}=0$ (19)
@@
$\neg s_{i}\wedge s_{\mathrm{q}}\neq 0\Rightarrow\{w_{1},w_{2}\}\not\subset W_{i}$ (20)
This binary test costs only a few machine instructions on data that is already in machine registers. It can thus be used as a pre-filter to before applying the predicate computation.
The remaining degree of freedom is how to choose the binary signatures, because this rule’s filtering ability depends on the choice of the signatures $S$. We experimented with iid Bernouilli bits with varying $p$.
@@
i.e. the best setting avoids to do the full check more than 4/5th of the times.
Pre- or post-filtering.
@@
There are two possible approaches to filtered search: post-filtering, which is described above, and pre-filtering, where only vectors with appropriate metadata are considered in vector search. The pre-filtering generates a subset of vectors to compare with that can then be compared in brute force. Brute force is slow but acceptable if the subset size is small.
@@
Therefore, the decision to use pre- or post-filtering depends on whether the subset of vectors is large or not. To this end we build a classical inverted list that maps each word $w$ to the list of items that contain $w$, of size $L_{w}$.
@@
If there is a single query word $\{w_{1}\}$ then the subset size is directly accessible as $L_{w_{1}}$. When there are two query words $\{w_{1},w_{2}\}$, finding the subset size requires intersecting the inverted lists for $w_{1}$ and $w_{2}$, which is slow. Instead, to take the decision, we estimate the subset size in the following way: we know that the empirical probability of each word to appear $P(w)=L_{w}/N$; assuming independent draws, the probability of both words to appear is $P(w_{1})P(w_{2})$. Therefore, the expected size of the intersection is $NP(w_{1})P(w_{2})=L_{w_{1}}L_{w_{2}}/N$.
We use this estimate to decide doing pre- or post-filtering. Of course, if pre-filtering is selected, the actual intersection still has to be computed.
7 Faiss engineering
@@
Faiss started in a research environment. As a consequence, it grew organically as indexing research was making progress. In the following, we briefly mention the guiding principles that kept the library coherent, how optimization is performed and an example of how Faiss internals are exposed so that it can be embedded in a vector database.
7.1 Code structure
@@
Faiss’s basic data types are concrete (not templates): vectors are always represented as 32-bit floats that are portable and provide a good tradeoff between size and accuracy. Similarly, all vector ids are represented with 64-bit integers. This is often larger than necessary for sequential numbering but is widely used for database identifiers.
7.2 High-level interface
@@
In order to facilitate experimentation, the whole library is wrapped for scripting languages such as Python with numpy. To this end, SWIG [^5] https://www.swig.org/ exhaustively generates wrappers for all C++ classes, methods and variables. The associated Python layer also contains benchmarking code, dataset definitions, driver code. More and more functionality is embedded in the contrib package of Faiss. Faiss also provides a pure C API, which is useful for bindings with programming languages such as Rust or Java.
@@ 7.3 Optimization
Often, only a subset of data sizes are optimized. For example, for PQ indexes, only $K=2^{8}$ and $K=2^{4}$ and $d/M\in\{2,4,8,16,20\}$ are fully optimized. For IndexLSH search, only code sizes 4, 8, 16, 20 are optimized. Fixing these sizes allows to write “kernels”, sequences of instructions without explicit loops or tests, that aim to maximize arithmetic throughput.
@@
Modern CPUs support Single Instruction, Multiple Data (SIMD) operations, specifically AVX/AVX2 for x86 and NEON for ARM. Faiss exploits those at three levels.
@@
The third level of optimization is to adapt the data layout and algorithms in order to speed up their SIMD implementation. The 4-bit product and additive quantizer implementations are implemented in this way, inspired by the SCANN library [Guo et al., 2020]: the layout of the PQ codes for several consecutive vectors is interleaved in memory so that a vector permutation can be used to perform the LUT lookups of (14) in parallel. This is implemented in the FastScan variants of PQ and AQ indexes (IndexPQFastScan, IndexIVFResidual QuantizerFastScan, etc.).
@@
Porting Faiss to the GPU is an involved undertaking due to substantial architectural specificities. The implementation of GPU Faiss is detailed in [Johnson et al., 2019], we summarize the GPU implementation challenges therein.
@@
Sequential table scanning in IVF indices requires loading data from main (global) memory. While latencies to access main memory are high, for table scanning we know in advance what data we wish to access, so the data movement from main memory into registers can be pipelined or use double buffering, so we can achieve close to peak possible performance.
Selecting the $k$ closest vectors to a query vector by ranking distances on the CPU is best implemented with a min- or max-heap. On the GPU, the sequential operations involved in heap operations would similarly force the GPU into a latency-bound regime. This is the largest challenge for GPU implementation of vector search, as the time needed for the heap implementation an order of magnitude greater than all other arithmetic. To handle this, we developed an efficient GPU $k$-selection algorithm [Johnson et al., 2019] that allows for ranking candidate vectors in a single pass, operating at a substantial fraction of peak possible performance per memory bandwidth limits. It relies upon heavy usage of the high-speed, large register memory on GPUs, and small-set bitonic sorting via warp shuffles with buffering techniques.
Irregular computations such as walking graph structures for graph-based indices like HNSW tend to remain in the latency-bound (due to the sequential traversal) rather than arithmetic throughput or memory bandwidth-bound regimes. Here, GPUs are at a disadvantage as compared to CPUs, and emerging techniques such as CAGRA [Ootomo et al., 2023] are required to parallelize otherwise sequential operations with graph traversal.
@@
Many Faiss components expose internal parameters to fine-tune the tradeoff between metrics: number of iterations of k-means, batch sizes for brute-force distance computations, etc. Default parameter values were chosen to work reasonably well in most cases.
@@
7.4 Interfacing with external storage
@@
This abstraction is useful when the inverted lists are not stored sequentially or fragmented into sub-lists because of metadata filtering [Huang et al., 2020]. Faiss is used only to perform the coarse quantization and the vector encoding.
8 Faiss applications
Faiss is used in many configurations. Hundreds of vector search applications rely on it, both within Meta and externally. Below we present a few use cases that showcase either an extreme scale or an application with particular impact.
8.1 Trillion scale index
For this example, Faiss is used to index 1.5 trillion vectors in 144 dimensions. The indexing needs to be accurate, therefore the compression of the vectors is limited to 54 bytes with a PCA to 72 dimensions and 6-bit scalar quantizer (PCAR72,SQ6).
Once the training is finished, the index is built by the following three phases:
  1. 1.
    shard over ids: add the input vectors in 2000 shards independently, producing 2000 indexes (each one fits in 256G RAM);
  2. 2.
  3. 3.
    load the shards: memory-map all 100 indices on a central machine as 100 OnDiskInvertedLists (a memory map of 83TiB).
Steps 1 and 2 are organized to be performed as independent cluster jobs on a few hundred servers (64 cores, 256G RAM). Otherwise, code uses standard Faiss in Python.
The central machine that handles searches performs the coarse quantization and loads the inverted lists from the distributed disk partition.
The limiting factor is the network bandwidth of this central machine. Therefore, it is more efficient to distribute the search on 20 intermediate servers to spread the load. This brings the search time down to roughly 1 s per query.
8.2 Text retrieval
Faiss is commonly used for natural language processing tasks. In particular, ANNS is relevant for information retrieval [Thakur et al., 2021, Petroni et al., 2021], with applications such as fact checking, entity linking, slot filling or open-domain question answering: these often rely on retrieving relevant content across a large-scale corpus. To that end, embedding models have been optimized for text retrieval [Izacard et al., 2021, Lin et al., 2023a].
Finally, [Izacard et al., 2023], [Lin et al., 2023b], [Shi et al., 2023b] and [Khandelwal et al., 2020] consist of language models that have been trained to integrate textual retrieval in order to improve their accuracy, factuality or compute efficiency.
8.3 Data mining
Another recurrent application of ANNS and Faiss is in the mining and curation of large datasets. In particular, Faiss has been used to mine bilingual texts across very large text datasets retrieved from the web [Schwenk and Douze, 2017, Barrault et al., 2023], or to organize a language model’s training corpus in order to group together series of documents covering similar topics [Shi et al., 2023a].
In the image domain, [Oquab et al., 2023] leverages Faiss to remove duplicates from a dataset containing 1.3B images. It then relies on efficient indexing in order to mine a curated dataset whose distribution matches the distribution of a target dataset.
8.4 Content Moderation
One of the major applications of Faiss is the detection and remediation of harmful content at scale. Human labelled examples of policy violating images and videos are embedded with models such as SSCD [Pizzi et al., 2022] and stored in a Faiss index. To decide if a new image or video would violate some policies, a multi-stage classification pipeline first embeds the content and searches the Faiss index for similar labelled examples, typically utilizing range queries. The results are aggregated and processed through additional machine classification or human verification. Since the impact of mistakes is high, good representations should discriminate perceptually similar and different content, and accurate similarity search is required even at billions to trillion-scale. The former problem motivated the Image and Video Similarity Challenges [Douze et al., 2021, Pizzi et al., 2023].
9 Conclusion
Over the years, Faiss has been expanding to include the most relevant vectors indexing techniques from research. We plan to continue doing this to include novel quantization techniques [Huijben et al., 2024], better hardware support for some types of indexes [Ootomo et al., 2023], and new forms of indexing, like associative vector memories for transformer architectures [Kang et al., 2024, Zhang et al., 2024].

### v3

@@ 6.2 Filtered search
This is sufficient to store enumerated types (e.g. country codes, music genres, license types, etc.), dates (as days since some origin), version numbers, etc. However, it is insufficient for more complex metadata. In the example use case below, we use the available bits to implement more complex filtering.
@@
In the filtered search track of the BigANN 2023 competition [75], each query and database vector is associated with a few terms from a fixed vocabulary of size $v$ (for the queries there are only 1 or 2 words). The filtering consists in considering only the database vectors that include all the query terms. . This metadata is given as a sparse matrix $M_{\mathrm{meta}}\in\{0,1\}^{N\times v}$.
@@
$\{w_{1},w_{2}\}\subset W_{i}\Rightarrow\neg s_{i}\wedge s_{\mathrm{q}}=0,$ (19)
@@
$\neg s_{i}\wedge s_{\mathrm{q}}\neq 0\Rightarrow\{w_{1},w_{2}\}\not\subset W_{i}.$ (20)
This binary test costs only a few machine instructions on data that is already in machine registers. It can thus be used as a pre-filter before applying the predicate computation. This is implemented in the module bow_id_selector [^6] https://github.com/harsha-simhadri/big-ann-benchmarks/tree/main/neurips23/filter/faiss.
The remaining degree of freedom is how to choose the binary signatures, because this rule’s filtering ability depends on the choice of the signatures $S$. We experimented with i.i.d. Bernoulli bits with varying $p$: the best setting avoids running the full predicate more than 4/5 times.
@@
Vector-first or metadata-first search.
@@
There are two possible approaches to filtered search: vector-first, which is described above, and metadata-first, where only vectors with appropriate metadata are considered in vector search. The metadata-first filtering generates a subset of vectors to compare with that can then be compared using brute force search. Brute force search is slow but acceptable if the subset size is small, and the results are exact.
@@
Therefore, in the context of the BigANN competition [75], the decision to use vector-first or metadata-first depends on how large the subset is. To this end, we map each word $w$ to the list of items that contain $w$, of size $L_{w}$.
@@
If there is a single query word $\{w_{1}\}$ then the subset size is directly accessible as $S=L_{w_{1}}$. With two query words $\{w_{1},w_{2}\}$, finding the subset size requires intersecting the inverted lists for $w_{1}$ and $w_{2}$, which is slow. Instead, one can estimate the size of the subset using the empirical probability of each word to appear $P(w)=L_{w}/N$. Assuming independent draws, the probability of both words to appear is $P(w_{1})P(w_{2})$. Therefore, the expected size of the intersection is $S\approx NP(w_{1})P(w_{2})=L_{w_{1}}L_{w_{2}}/N$. If $S/N<3\times 10^{-4}$ (an empirical threshold), then the subset is sufficiently small that metadata-first can be applied. Of course, if metadata-first is selected, the actual intersection has to be computed. Most participants to the competition used similar heuristics.
7 Faiss applications
Faiss is widely used across the industry, with numerous applications leveraging its capabilities. The following examples highlight notable use cases that demonstrate exceptional scalability or significant impact.
7.1 Trillion scale index
In this example, we index 1.5 trillion vectors in 144 dimensions. The indexing needs to be accurate, therefore the compression of the vectors is limited to 54 bytes with a PCA to 72 dimensions and 6-bit scalar quantizer (PCAR72,SQ6).
Once the training is completed, the index is built in the following three phases:
  1. 1.
    shard over ids: add the input vectors in 2000 shards independently, producing 2000 indexes (each one fits in 256 GB RAM);
  2. 2.
  3. 3.
    load the shards: memory-map all 100 indexes on a central machine as 100 OnDiskInvertedLists (a memory map of 83 TiB).
Steps 1 and 2 are organized to be performed as independent cluster jobs on a few hundred servers (64 cores, 256G RAM). Otherwise, the code is written in standard Faiss in Python.
The central machine that handles searches performs the coarse quantization and loads the inverted lists from the distributed disk partition. The limiting factor is the network bandwidth of this central machine. Therefore, it is more efficient to distribute the search on 20 intermediate servers to spread the load. This brings the search time down to roughly 1 s per query.
7.2 Text retrieval
Faiss is commonly used for knowledge intensive natural language processing tasks. In particular, ANNS is relevant for information retrieval [85, 67], with applications such as fact checking, entity linking, slot filling or open-domain question answering: these often rely on retrieving relevant content across a large-scale corpus. To that end, embedding models have been optimized for text retrieval [40, 49].
Finally, [41], [50], [74] and [48] consist of language models that have been trained to integrate textual retrieval in order to improve their accuracy, factuality or compute efficiency.
7.3 Data mining
Another recurrent application of ANNS and Faiss is in the mining and curation of large datasets. In particular, Faiss has been used to mine bilingual texts across very large text datasets retrieved from the web [72, 9], or to organize a language model’s training corpus in order to group together series of documents covering similar topics [73].
In the image domain, [64] leverages Faiss to remove duplicates from a dataset containing 1.3B images. It then relies on efficient indexing in order to mine a curated dataset whose distribution matches the distribution of a target dataset.
7.4 Content Moderation
One of the major applications of Faiss is the detection and remediation of harmful content at scale. Human-labeled examples of policies violating images and videos are embedded with models such as SSCD [69] and stored in a Faiss index. To decide if a new image or video would violate some policies, a multi-stage classification pipeline first embeds the content and searches the Faiss index for similar labeled examples, typically utilizing range queries. The results are aggregated and processed through additional machine classification or human verification. Since the impact of mistakes is high, good representations should discriminate perceptually similar and different content, and accurate similarity search is required even at billion to trillion scale. The former problem motivated the Image and Video Similarity Challenges [30, 68].
8 Conclusion
Throughout the years, Faiss continuously expanded its focus to include the most relevant vector indexing techniques from research. We plan to continue doing this to include novel quantization techniques [39], better hardware support for some indexes [63] and new indexing forms, such as associative vector memories for transformer architectures [20, 92].
Appendix A Appendix
This appendix exposes aspects of Faiss’s implementation. Faiss started in a research environment. As a consequence, it grew organically as indexing research was making progress.
@@ Appendix A Appendix
In the following, we summarize the guiding principles that keep the library coherent (Appendix A.1); the structure of the library and its dependencies (Appendix A.2); how optimization is performed (Appendix A.3); an example of how Faiss internals are exposed so that it can be embedded in a vector database (Appendix A.4; and finally, a flowchart that shows how to choose a Faiss index (Appendix A.5).
A.1 Code structure
@@
Faiss’s basic data types are concrete (not templates): vectors are always represented as 32-bit floats that are portable and provide a good trade-off between size and accuracy. Similarly, all vector ids are represented with 64-bit integers. This is often larger than necessary for sequential numbering but is widely used for database identifiers.
Faiss is modular and includes few dependencies, so that linking it from C++ is easy. On the Python level, callback classes (ResultHanlder, InvertedLists, IDSelector) can be subclassed and wrapped in SWIG so that they can be provided to Faiss without recompiling Faiss itself.
A.2 High-level interface
@@
In order to facilitate experimentation, the whole library is wrapped for Python with numpy. To this end, SWIG [^7] https://www.swig.org/ exhaustively generates wrappers for all C++ classes, methods and variables. The associated Python layer also contains benchmarking code, dataset definitions, driver code. More and more functionality is embedded in the contrib package of Faiss. Faiss also provides a pure C API, which is useful for bindings with programming languages such as Rust or Java.
@@ A.3 Optimization
Often, only a subset of data sizes are optimized. For example, for PQ indexes, only $K$ = $2^{8}$ and $K$ = $2^{4}$ and $d/M\in\{2,4,8,16,20\}$ are fully optimized. For IndexLSH search, we only optimized code sizes 4, 8, 16 and 20. Fixing these sizes allows to write dedicated “kernels”, i.e., sequences of instructions without explicit loops or tests, that aim to maximize arithmetic throughput.
@@
Modern CPUs support Single Instruction, Multiple Data (SIMD) operations, specifically AVX/AVX2/AVX512 for x86 and NEON for ARM. Faiss exploits those at three levels.
@@
The third level of optimization adapts the data layout and algorithms in order to speed up their SIMD implementation. The 4-bit product and additive quantizer implementations are implemented in this way, inspired by the SCANN library [36]: the layout of the PQ codes for several consecutive vectors is interleaved in memory so that a vector permutation can be used to perform the LUT lookups of (14) in parallel. This is implemented in the FastScan variants of PQ and AQ indexes (IndexPQFastScan, IndexIVFResidual QuantizerFastScan, etc.).
@@
Porting Faiss to the GPU is an involved undertaking due to substantial architectural specificities. The implementation of GPU Faiss is detailed in [47], we summarize the GPU implementation challenges therein.
@@
Sequential table scanning in IVF indices requires loading data from main (global) memory. While main memory access latency is high, for table scanning we know in advance what data we wish to access. The data movement from main memory into registers can thus be pipelined or use double buffering, so we can achieve close to peak possible performance.
Selecting the $k$ closest vectors to a query vector by ranking distances on the CPU is best implemented with a min- or max-heap. On the GPU, the sequential operations involved in heap operations would similarly force the GPU into a latency-bound regime. This is the largest challenge for GPU implementation of vector search, as the time needed for the heap implementation an order of magnitude greater than all other arithmetic. To handle this, we developed an efficient GPU $k$-selection algorithm [47] that allows for ranking candidate vectors in a single pass, operating at a substantial fraction of peak possible performance per memory bandwidth limits. It relies upon heavy usage of the high-speed, large register memory on GPUs, and small-set bitonic sorting via warp shuffles with buffering techniques.
Irregular computations such as walking graph structures for graph-based indices like HNSW tend to remain in the latency-bound (due to the sequential traversal) rather than arithmetic throughput or memory bandwidth-bound regimes. Here, GPUs are at a disadvantage as compared to CPUs, and emerging techniques such as CAGRA [63] are required to parallelize otherwise sequential operations with graph traversal.
@@
Many Faiss components expose internal parameters to fine-tune the trade-off between metrics: number of iterations of k-means, batch sizes for brute-force distance computations, etc. Default parameter values are set to work reasonably well in most cases.
@@
A.4 Interfacing with external storage
@@
This abstraction is useful when the inverted lists are not stored sequentially or fragmented into sub-lists because of metadata filtering [38]. Faiss is used only to perform the coarse quantization and the vector encoding.
A.5 How to choose an index
Figure 10 shows a decision tree for index types, depending on database size and memory constraints. The decision tree first focuses on the “hard” memory constraint, then on secondary trade-offs, like index construction time vs. accuracy.
Figure 10: Decision tree to choose a Faiss index. This is for the common case of Euclidean k-nearest neighbor search on CPU. The resulting indexes (in bold) are defined by their factory string (see Appendix A.2).
A.6 API index of the Faiss library
As mentioned in Section 5.3, most indexes are a combination of a compression method (like PQ) and a non-exhaustive search method. The index class names are built as shown in Table 3.
Figure 11 shows the most important index classes in Faiss, grouped by broad families. Some combinations are omitted for brevity.
No encoding PQ encoding scalar quantizer
    Flat IndexFlat IndexPQ IndexScalarQuantizer
    IVF IndexIVFFlat IndexIVFPQ IndexIVFScalarQuantizer
    HNSW IndexHSNWFlat IndexHNSWPQ IndexHNSWScalarQuantizer
Faiss index API.
In the following, we list the main methods offered by Faiss indexes, $x\in\mathbb{R}^{n\times d}$ is a list of $n$ vectors in dimension $d$ represented as the rows of a matrix, and $I\in\{0,2^{63}-1\}^{n}$ is a list of 63-bit ids of size $n$.
  • train($x$) perform a training using vectors $x$ to prepare adding vectors of the same data distribution to the index;
  • add($x$) add the vectors $x$ to the index, numbered sequentially;
  • add_with_ids($x$,$I$) add the vectors $x$, identified by the 63-bit ids $I$;
  • search($x$, $k$) return the $k$ nearest vectors of each of the query vectors in $x$;
  • range_search($x$, $\varepsilon$) return all vectors within a radius $\varepsilon$ of each of the query vectors in $x$;
  • remove_ids($I$) remove vectors with ids $I$ from the index;
  • reconstruct_batch($I$) extract the vectors with ids in $I$. It is called “reconstruct” because for most index types, the returned vectors will be approximations of the original ones.
Figure 11: Hierarchy of the main index classes in CPU Faiss for floating-point vectors. For each class we indicate the corresponding factory string (when applicable) and a short explanation.
Quantizer objects.
Table 4 shows the hierarchy of quantizers. Each quantizer can represent the reproduction values of all the quantizers below, but is slower to train and to perform assignment with.
The root Quantizer class has the following fields and methods (where $x\in\mathbb{R}^{n\times d}$ is a list of vectors):
  • code_size size of the codes it produces (in bytes);
  • train($x$) perform a training using vectors $x$ to prepare the quantizer;
  • compute_codes($x$) encode a $n$ vectors to an array of size $n\times$(code_size);
  • decode($C$) decode the vectors from codes obtained with encoder;
Table 4: The hierarchy of quantizers. Each quantizer can represent the set of reproduction values of the quantizers below it.
Type of quantizer class
Vector quantizer
Additive quantizer AdditiveQuantizer
Product-additive quantizer ProductAdditiveQuantizer
Product quantizer ProductQuantizer
Scalar quantizer ScalarQuantizer
Binarization